Report of Activities on the RVIB N.B. Palmer
11 August 2001

The work of the broad-scale survey has been temporarily stopped to carry out a planned mid-cruise rendezvous with the R/V L.M. Gould. What was not planned was the need for us to steam all the way to their current position, which was about 50 nm to the northeast from station #51, the last station we completed before starting the steam. They were planning to meet us halfway, but found that they were stuck in the ice floe that they had been working in for the past few days or so and could not move. So we are on the way to break them out of their current position, help them get underway to their new work site, and to do a transfer of some equipment and a person to the Palmer. One item coming to the Palmer is an intermediate ice buoy, which the Gould had planned to deploy on the southern end of the survey grid. But the logistics of getting to the location became too difficult and so this task will be undertaken by the Palmer. An item going from the Palmer to the Gould is a band saw for cutting up ice cores. The Gould's band saw self-destructed and could not be repaired. We are currently (1000 on 13 August) at -68° 07.593S; -70° 16.662 W. Winds are 10 to 15 kts out of the west (270) and the air temperature is -16.2°C. The skies are partly cloudy and there is morning sunshine- a welcome change. The barometer is still low at 959.2 mlb, but with the wind shift and the precipitous drop in temperature, it appears the low is now to our east and higher pressure should be building.

On 11 August, work was completed at only two stations, #47 and 48. Station #47 took an especially long time because it was a full deep station at one of the furthest points offshore in the survey grid. Snow was again with us most of the day along with moderate temperatures around freezing and winds out of the northwest in the 20 to 30 kt range. At station #47, there were patches of pack ice that, when over-turned in our wake after being pushed aside by the ship, had a light tan to dark brown appearance, a sign that they had considerable ice algae growing on them and that these pieces of ice were older. We have not seen many of these heavily colonized pieces of ice in the more inshore portions of our survey area. Work completed at these two stations included two CTD casts to the bottom, a 1-m2 MOCNESS to 1000 m, and a phytoplankton ring net tow an ROV under ice survey, and a Tucker trawl. BIOMAPER-II was returned to service and towyoed between stations #47, 48, and 49.

John Klinck reports that on 11 August, the CTD group did two stations at the offshore end of line 260.

Station #47 (2927 m) is in oceanic water off the shelf break. The mixed layer is uniform in character to a depth of about 85 m. There is no indication of layering in the pycnocline. The O2 minimum sits at about 280 m. The temperature maximum is split by a colder layer producing a double temperature maximum at 280 and 380 m. Water below 500 m is typical of this region with nearly constant (and slightly decreasing) salinity, decreasing temperature and increasing O2 with depth.

Station #48 (390 m) is on the outer shelf, but is typical of shelf stations. The mixed layer is uniform to 80 m and joins a ragged pycnocline with strong and common layers. The water below the pycnocline also shows evidence of strong layers with numerous temperature reversals with layers reaching 10 m thickness.

Ari Friedlaender reports that on 11 August, marine mammal observations began at 1200, position 67° 12.42S, 74° 34.43W. Ice conditions were 10/10 of vast unbroken floes with low ridging. There were sparse pools of open water every once in a while, but snow fog limited our visibility to less that 0.5 nm, with realistic sight-ability even less than that. Effort continued until light conditions diminished shortly after 1600. No cetaceans were seen.

Ana Sirovic reports that on 11 August, she heard seals on the sonobuoy that was deployed on 11 August! There were at least 2 (possibly 3) animals calling over a 12 min interval. She believes they were Weddell seals. This buoy was deployed when we were in the vicinity of one of the moored instruments (S6 at 67° 17.9, 74° 10.8), in transit between stations #47 and 48. The buoy was monitored for a total of 1 h 23 min, before it was out of the antenna range.

BIOMAPER-II/MOCNESS report (P. Wiebe, C. Ashjian, and S. Gallager):
The problems with the ESS sensor system, which also serves as the controller for the power to the Video Plankton Recorder on BIOMAPER-II were finally solved. A short test during the wee hours at station #47 demonstrated that it would work in seawater. It was deployed for the steam between stations #47 and 48 around 1300 on 11 August. Volume backscattering along this transit was very low in the upper 300 m, although a layer was evident between 75 and 100 m in which pteropods were seen in the VPR images. After leaving station #48, a relatively dense layer of scatterers were observed near the bottom that extended to about 100 meters above the bottom. Bottom depths were around 400 m. This feature was observed on other transect lines as we traveled across the outer portion of the continental shelf. We have not yet, however, been able to get a net down into the layer to sample the organisms present there and so we are not able to infer what species might be responsible for the increased volume backscattering. Noise, whether ambient or instrumental, is always an issue with high frequency acoustical systems. Large amounts of noise have occurred in the down looking 43 and 120 kHz echograms and there has been some suspicion that it might be instrumental. Thus, coming into station both stations #48 and 49, noise tests were conducted. One test indicated that our noise levels were lower than those observed in the Magellan Straits at the start of the cruise. The other clearly showed that a large fraction of what we have been recording when the ship is using four engines to drive the ship through heavy pack ice is coming from noise generated by the ship's machinery, propeller turbulence, and the ice.

MOCNESS tow #11 was conducted in deep water off of the shelf break at station #47. The tow went to 1000 m and was quite successful, except for difficulties with cross-talk between the OPC and the MOCNESS. This resulted in the OPC being turned off and not sampling the upper 100 m.

Copepods were abundant from 100-1000 m. Three Cyclothone and one large myctophid were collected from 800-1000 m. Ostracods and chaetognaths were found in the 600-800 m range, along with two large myctophids and many jelly animals, which were quite mangled and not easy to identify. Krill (Euphausia and Thysanoessa) were found from 400-600 m. Chaetognaths and jelly animals were found from 200-400 m. Many Thysanoessa were seen from 100-200 m, along with chaetognaths and ostracods. The upper water column had very low abundances. Euphuasiids and ostracods were observed from 50-100 m. The upper 50 m had very little, with only a few copepods and amphipods and a fish larva in the 25-50 m range. The upper 25 m collected a significant quantity of ice.

Cheers, Peter